OpenStudy (anonymous):

find the derivative of f(x)=√x +1

OpenStudy (luigi0210):

Try the power rule?

OpenStudy (luigi0210):

Wait, it is \(\Large f(x)=\sqrt{x+1}\) or \(\Large f(x)=\sqrt{x}+1\)?

OpenStudy (anonymous):

the second one

rvc (rvc):

d(x^n)/dx =n(x)^n-1

OpenStudy (luigi0210):

Yea, use the power rule, take it in this form: \(\Large f(x)=(x)^{1/2}+1\)

OpenStudy (michele_laino):

then: \[f'(x)=\frac{ 1 }{ 2\sqrt{x} }\]

OpenStudy (anonymous):

how come square root x stays at the bottom ?

OpenStudy (luigi0210):

The power turn negative.

OpenStudy (luigi0210):

You bring down the power, then subtract one.

OpenStudy (midhun.madhu1987):